• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 08:00:47 on July 9, 1999

    Entry number 22066

    keyword=Beam Time Accounting
    Hall A

    7/9/1999 0000-0800
    Experiment #: E91-010
    Shift Leader: Humensky
    Shift Workers: Younus

    00000100020003000400050006000700Total Time
    Acceptable Beam
    in Use
    0 0 0 5 30 1 40 40 2.91666666666667
    Beam Available
    but Not in Use
    0 0 0 0 0 0 0 0 0
    Beam Not Available
    or Unacceptable
    1 1 1 55 0 0 5 0 4
    Accelerator Con-
    figuration Chage
    0 0 0 0 30 0 15 20 1.08333333333333
    Experiment
    Ready
    1 1 1 1 1 1 1 1 8
    Planned Config-
    uration Change
    0 0 0 0 0 0 0 0 0
    Unplanned
    Experiment Down
    0 0 0 0 0 0 0 00
    Hall not
    ready reasons
    Target 0 0 0 0 0 0 0 0 0
    HRSe 0 0 0 0 0 0 0 0 0
    HRSh 0 0 0 0 0 0 0 0 0
    Data Acq 0 0 0 0 0 0 0 0 0
    Other 0 0 0 0 0 0 0 0 0

    Comments: Beam was taken away to tune to Hall B at 23:20; beam returned for the Hall A current calibration at 00:25. 00:35: Saha just spoke with the PD--MCC had accidentally loaded the Moller quad settings rather than the regular Hall A achromatic setting. This explains why the tune was too dirty for an accurate Arc energy measurement. In resetting the magnets MCC located some of their other tune troubles... so at 01:00 the update was 1-1.5 hours down. Neyret made an opportunistic access for Compton during that time. About 02:30 we got a little pulsed beam and tripped a Moller magnet. As of 03:20 MCC has completed their access to reset the magnet and is in the process of bringing it up. CW beam 3uA at 03:55. 05:00: Arc energy measurement finished; setting up for ep energy measurement. 06:10: Finished ep measurement. Starting current calibration. Reference DVM in injector is not stable so couldn't complete measurement. Did a little trouble-shooting of bcm calibration equipment and did a sweet spot scan.

    Data Entry by: Humensky